### v4

@@ 2.3 Causal Estimands
Because we allow for cross-cluster interference, we need to redefine conventional estimands in a manner free of this source of bias. To this end, define for $t\in\{0,1\}$ the $p_{t}$-counterfactual design, which sets $k=1$ and $q=t$ in Assumption 4. This design groups the population into a single cluster and assigns treatment with probability $p_{t}$. Throughout the paper, let ${\bf E}[\cdot]$ denote the expectation taken with respect to the observed design in Assumption 4 and ${\bf E}^{*}_{p_{t}}[\cdot]$ the expectation taken with respect to the $p_{t}$-counterfactual design.
@@ 5.1 Empirical Examples
Sur et al. (2009) conduct a CRT in an urban slum in India spanning about 1.2 by 0.7 km with 38k participants. To compute our suggested number of clusters (9), we need to select $\tilde{\gamma}$ and the unit of length. Suppose we choose the conservative bound $\tilde{\gamma}=d=2$, so that interference is assumed to decay like $r^{-2}$ or faster with each unit of length $r$. If we take 35m to be the unit of length, then for units $i,j,k$ such that $\rho(i,j)=35$m and $\rho(i,k)=70$m, the extent to which $k$’s treatment affects $i$ is less than $1/4$th ($0.252^{-\tilde{\gamma}}=$ 0.25) as small as the extent to which $j$’s affects $i$. This is with only a 35m difference in distance. For this unit of length, (9) yields $k=78$. In other words, these are the assumptions on interference that justify the authors’ choice of $k=80$, originally determined by a conventional power analysis based on a partial interference model.
@@ 6 Simulation Study
We conduct a simulation study to illustrate the finite-sample properties of our estimator and difference in means under our proposed design. We randomly draw unit locations from the square $[-(n\alpha_{n})^{1/2},(n\alpha_{n})^{1/2}]^{2}$ with $\alpha_{n}=0.8,0.7,0.6$, respectively. This corresponds to the infill-increasing case where the regional volume shrinks with the population size. We create clusters using $k$-medoids with $k$ given by (9) using the conservative choice $\tilde{\gamma}=d=2$ and set the assignment probabilities in Assumption 4 to $(q,p_{1},p_{0})=(0.7,0.5,0)$.
@@
Due to the choice of $-5$ in the spatial weights $w_{ij}$, Assumption 3 holds for $\gamma=3$ (Leung, 2022b, eq. (3)), which is a fairly slow rate of decay given that Assumption 3 requires $\gamma>2$.
@@
We present results for the indirect and total effects using 5000 simulation draws where within each draw, we redraw potential outcomes and recompute the design-based estimand. In Table 1, the “Spatial Interference” columns correspond to the outcome model described above, whereas the “Partial Interference” columns redefine $w_{ij}=0$ if $i,j$ lie in different clusters to eliminate cross-cluster interference. The “CI” rows report the coverage of 95-percent CIs using the indicated standard errors. For our estimator, the “SE” row corresponds to standard errors obtained from our variance estimator (4), while for difference in means, it corresponds to conventional cluster-robust SEs. The “SE” rows are the true superpopulation standard errors obtained by taking the standard deviation of the estimator across the simulation draws. As such, SE should be consistent for SE rather than conservative, while the CIs have asymptotically conservative coverage for the design-based estimand. Finally, row “Excl” is the percentage of units that are not well surrounded.
@@ 7 Empirical Application
Egger et al. (2022) estimate the effects of the transfers on the population of eligible households using two main specifications described in their §3.2. Their “RF” (reduced form) specification is an OLS regression of an outcome on household-level and sublocation-level treatment indicators and covariates. They report the coefficient on the household-level indicator, which corresponds to a model-assisted estimate of the direct effect $\theta^{*}(1,0,2/3,2/3)$ for eligible households.
However, the authors note the potential for interference across sublocations (their quote in Example 2). For this reason, their preferred specification is the following “IV” (instrumental variables) regression. The main regressors are the amount of cash transferred per capita to the household’s village and the amount transferred to neighboring villages within a band of $r-2$ to $r$ km from the ego’s village for a range of $r$ values. The corresponding instruments are respectively an indicator for the ego’s village being treated and the share of eligible households in the band assigned to treatment. They use a BIC criterion to select the maximum range of $r$, which is 2 km. Using these estimates, the authors compute a model-assisted estimate of the total effect $\theta^{*}(1,0,2/3,0)$. This compares saturation levels of two-thirds and zero, which is nonparametrically unidentified since control villages have a saturation level of one-third. We instead report our estimates of the total effect $\theta^{*}(1,0,2/3,1/3)$.
@@ 8 Conclusion
    Initialize an arbitrary set of $k$ medoids =$\mathcal{M}=\{m_{j}\}_{j=1}^{k}\subseteq\mathcal{N}_{n}$. Given any $\mathcal{M}$, the associated set of clusters is $\mathcal{C}_{n}=\{C_{j}\}_{j=1}^{k}$, where $C_{j}=\{i\in\mathcal{N}_{n}\colon m_{j}=\operatornamewithlimits{argmin}_{m\in\mathcal{M}}\rho(i,m)\}$.
@@
    While there exist $m\in\mathcal{M}$ and $o\in\mathcal{N}_{n}\backslash\mathcal{M}$ such that $\text{cost}(\mathcal{M}\backslash\{m\}\cup\{o\})<\text{cost}(\mathcal{M})$, replace $m$ with the $o$ that minimizes $\text{cost}(\mathcal{M}\backslash\{m\}\cup\{o\})$.
@@
It would also be useful to develop designs for nonparametrically estimating $\gamma$. A preliminary proposal is the following. Given $k$ clusters, say constructed using $k$-medoids, denote by $m_{j}$ the centroid of a cluster $C_{j}$. Randomize clusters to the following $T$ treatment arms. In the 0th arm, we assign all units to control. In the $t$th treatment arm for $t\geq 1$, we only assign units $i$ for which $\rho(i,m)\in(t,t+1]$ to treatment, where $m$ is the centroid of the cluster in question. That is, we treat only units in a ring at a certain distance from the centroid. Let $\hat{\theta}_{t}$ denote the difference-in-means estimate comparing average outcomes of units near centroids of clusters assigned to arm $t\geq 1$ with those assigned to arm 0. Then under Assumption 3, $\hat{\theta}_{t}$ should decay like $t^{-\gamma}$, producing a sort of “causal covariogram.” We may then regress $\log\hat{\theta}_{t}$ on $\log t$ to estimate $-\gamma$.
@@
  • $653$ villages (units), $84$ sublocations (clusters). Standard errors are in parentheses. Column IV is the IV estimate of the overall effect from Table I, column 3 of Egger et al. (2022).
@@
For any $i\in\mathcal{N}_{n}$, $d_{t}\in\{0,1\}$, $\bm{d}\in\{0,1\}^{n}$, and $S\subseteq\mathcal{N}_{n}$ containing $i$, we use the notation $(d_{t},\bm{D}_{S\backslash\{i\}},\bm{d}_{-S})$ to mean that we take the observed treatment vector $\bm{D}$, replace entry $D_{i}$ with $d_{t}$, and replace the subvector $\bm{D}_{\mathcal{N}_{n}\backslash S}=(D_{j}\colon j\in\mathcal{N}_{n}\backslash S)$ with the corresponding entries of $\bm{d}_{\mathcal{N}_{n}\backslash S}$.
@@
By Lemma SA.5.4, $\max_{i}\lvert\Lambda_{i}\rvert\precsim n/k_{n}$, so $[P1]\precsim k_{n}^{-1}$. It remains to show that $[P2]\precsim k_{n}^{-1}$. Intuitively, $[P2]$ should be well controlled since units not in $\Lambda_{i}$ are “far” from and therefore less correlated with $i$.
@@
We next show that (SA.6.9)$\precsim\xi_{n}/n$. By Assumptions 2(a) and 5, $\max_{\ell}\lvert C_{\ell}\rvert\precsim n/k_{n}$. By Assumption 2(b), $\sum_{j\not\in\Lambda_{i}}\bm{1}\{\rho(i,j)\in[s,s+1)\}\leq\lvert\mathcal{N}(i,s+1)\backslash\mathcal{N}(i,s)\rvert\leq C\xi_{n}\max\{s^{d-1},1\}$. Then
$[P2.1]=c_{n}^{\prime}\frac{k_{n}}{n^{2}}\frac{n}{k_{n}}\sum_{s=\lfloor 2r_{n}\rfloor}^{\lceil 4\bar{R}\rceil}C\xi_{n}s^{d-1}r_{n}^{-\gamma}\precsim\frac{\xi_{n}}{n}\bar{R}^{d}r_{n}^{-\gamma}$ (SA.6.10)
@@
(SA.6.10)$\precsim\frac{\xi_{n}}{n}\left(\frac{n}{k_{n}\xi_{n}}\right)^{1-\gamma/d}=\frac{1}{k_{n}}\left(\frac{k_{n}\xi_{n}}{n}\right)^{\gamma/d}\precsim\frac{\xi_{n}}{n}$ (SA.6.12)
@@
$[P2.2]=c_{n}^{\prime}\frac{k_{n}}{n^{2}}\frac{n}{k_{n}}\sum_{s=\lfloor 4\bar{R}\rfloor}^{\infty}C\xi_{n}s^{d-1}(s/2-\bar{R})^{-\gamma}\\ \leq Cc_{n}^{\prime}\frac{\xi_{n}}{n}\sum_{s=\lfloor 4\bar{R}\rfloor}^{\infty}s^{d-1}(s/4)^{-\gamma}\precsim\frac{\xi_{n}}{n}\sum_{s=1}^{\infty}s^{d-1-\gamma}.$ (SA.6.13)
Combining (SA.6.9), (SA.6.10), (SA.6.12), and (SA.6.13) yields $[P2]\precsim\xi_{n}/n\precsim k_{n}^{-1}$ since $k_{n}\xi_{n}/n\precsim 1$ by assumption.
@@
$\lvert{\bf E}[\tilde{\theta}^{+}]-\theta^{*}\rvert=\bigg{|}\frac{1}{n}\sum_{i\in\mathcal{N}_{n}}\big{(}{\bf E}[Y_{i}\mid T_{1i}^{+}=1]-{\bf E}^{*}_{p_{1}}[Y_{i}(d_{1},\bm{D}_{-i})]\big{)}\bigg{|}\\ =\beta\frac{1}{n}\sum_{i\in\mathcal{N}_{n}}\frac{\sum_{j\in\mathcal{N}_{n}\backslash C_{c(i)}}\bm{1}\{\rho(i,j)\leq 2\}}{\sum_{j\in\mathcal{N}_{n}}\bm{1}\{\rho(i,j)\leq 2\}}.$ (SA.6.18)
@@
~$\tilde{\mathcal{B}}(C)=\big{\{}i\in C\colon\lvert\mathcal{N}(i,1)\cap(\{\xi_{n}^{-1/d}x\colon x\in\mathbb{Z}^{d}\}\backslash C)\rvert\geq 1\big{\}},$
@@
(SA.6.18)$\geq\beta\frac{1}{n}\sum_{j=1}^{k_{n}}\sum_{i\in\tilde{\mathcal{B}}(C_{j})\backslash\tilde{\mathcal{B}}(\mathcal{N}_{n})}\frac{\sum_{\ell\in\mathcal{N}_{n}\backslash C_{j}}\bm{1}\{\rho(i,\ell)\leq 2\}}{\sum_{\ell\in\mathcal{N}_{n}}\bm{1}\{\rho(i,\ell)\leq 2\}}.$ (SA.6.19)
@@
$\min_{j=1,\ldots,k_{n}}\min_{i\in\tilde{\mathcal{B}}(C_{j})\backslash\tilde{\mathcal{B}}(\mathcal{N}_{n})}\frac{\sum_{\ell\in\mathcal{N}_{n}\backslash C_{j}}\bm{1}\{\rho(i,\ell)\leq 2\}}{\sum_{\ell\in\mathcal{N}_{n}}\bm{1}\{\rho(i,\ell)\leq 2\}}\succsim 1.$
@@
(SA.6.19)$\succsim\frac{1}{n}\bigg{(}\sum_{j=1}^{k_{n}}\lvert\tilde{\mathcal{B}}(C_{j})\rvert-\lvert\tilde{\mathcal{B}}(\mathcal{N}_{n})\rvert\bigg{)}.$ ) . (SA.6.20)
@@
|~$\lvert\tilde{\mathcal{B}}(\mathcal{N}_{n})\rvert\sim\xi_{n}\mathcal{R}_{n}^{d-1}\sim n(\xi_{n}/n)^{1/d}$ (SA.6.21)
Let $m_{j}$ denote the medoid of a cluster $C_{j}$. By Assumption 5(a), units in ~$\tilde{\mathcal{B}}(C_{j})$ are at least distance $\lfloor L_{n}\rfloor$ from $m_{j}$. Define the “contour set”
@@
|~$\lvert\tilde{\mathcal{B}}(C_{j})\rvert\geq\lvert\tilde{J}(\lfloor L_{n}\rfloor,C_{j})\rvert\sim\xi_{n}(n/(k_{n}\xi_{n}))^{(d-1)/d}$
@@
$\text{Var}\left(\frac{1}{n}\sum_{i\in\mathcal{N}_{n}}Z_{i}^{+}\right)\geq\alpha\frac{1}{n^{2}}\sum_{i\in\mathcal{N}_{n}}\sum_{j\in C_{c(i)}}\sum_{\ell\in C_{c(i)}\backslash\{i\}}\frac{\bm{1}\{\rho(i,\ell)\leq 2\}}{\sum_{\ell^{\prime}\in\mathcal{N}_{n}}\bm{1}\{\rho(i,\ell^{\prime})\leq 2\}}\\ \times\sum_{m\in C_{c(i)}\backslash\{j\}}\frac{\bm{1}\{\rho(j,m)\leq 2\}}{\sum_{m^{\prime}\in\mathcal{N}_{n}}\bm{1}\{\rho(j,m^{\prime})\leq 2\}}\succsim\frac{1}{n^{2}}\sum_{i\in\mathcal{N}_{n}}\lvert C_{c(i)}\rvert.$
@@
The right-hand side equals (SA.6.9) multiplied by $k_{n}$. Since (SA.6.9)$\precsim\xi_{n}/n$ as shown in the proof of Theorem 4, this is $\precsim k_{n}\xi_{n}/n$, as desired.
@@
which is $\prec 1$ by the proof of Theorem 2. Hence, (SA.6.29)1precedesitalic-(SA.6.29italic-)$\prec 1$.
@@
(SA.6.30)$\precsim\bigg{(}\frac{1}{n^{2}}\sum_{i\in\mathcal{N}_{n}}\sum_{j\in\Lambda_{i}}\lvert\text{Cov}(Z_{i},Z_{j})\rvert+\frac{1}{n^{2}}\sum_{i\in\mathcal{N}_{n}}\sum_{j\not\in\Lambda_{i}}\lvert\text{Cov}(Z_{i},Z_{j})\rvert\bigg{)}^{1/2}.$
@@
$\frac{k_{n}}{n^{2}}\sum_{i\in\mathcal{N}_{n}}\sum_{j\in\mathcal{N}_{n}}(Z_{i}-{\bf E}[Z_{i}])&(Z_{j}-{\bf E}[Z_{j}])A_{ij}(1)\\ &-\frac{k_{n}}{n^{2}}\sum_{i\in\mathcal{N}_{n}}\sum_{j\in\mathcal{N}_{n}}\text{Cov}({\bf E}[Z_{i}\mid\mathcal{F}_{i}],{\bf E}[Z_{j}\mid\mathcal{F}_{j}])A_{ij}(1)\end{aligned}$
[…]
@@
As previously argued, $\max_{i}\sum_{j\in\mathcal{N}_{n}}{\bf E}[Z_{j}]A_{ij}(1)\precsim n/k_{n}$, and $\lvert n^{-1}\sum_{i\in\mathcal{N}_{n}}(Z_{i}-{\bf E}[Z_{i}])\rvert\prec 1$ by the proof of Theorem 2, so $[P2]\prec 1$, while
$[P1]=\frac{k_{n}}{n^{2}}\sum_{i\in\mathcal{N}_{n}}\sum_{j\in\mathcal{N}_{n}}(Z_{i}-{\bf E}[Z_{i}\mid\mathcal{F}_{i}])(Z_{j}-{\bf E}[Z_{j}\mid\mathcal{F}_{j}])A_{ij}(1)\\ +2\frac{k_{n}}{n^{2}}\sum_{i\in\mathcal{N}_{n}}\sum_{j\in\mathcal{N}_{n}}{\bf E}[Z_{i}\mid\mathcal{F}_{i}](Z_{j}-{\bf E}[Z_{j}\mid\mathcal{F}_{j}])A_{ij}(1)\\ +\frac{k_{n}}{n^{2}}\sum_{i\in\mathcal{N}_{n}}\sum_{j\in\mathcal{N}_{n}}({\bf E}[Z_{i}\mid\mathcal{F}_{i}]{\bf E}[Z_{j}\mid\mathcal{F}_{j}]-{\bf E}[{\bf E}[Z_{i}\mid\mathcal{F}_{i}]{\bf E}[Z_{j}\mid\mathcal{F}_{j}]])A_{ij}(1)\\ \equiv[P1.1]+[P1.2]+[P1.3].$
@@
$[P1.1]\precsim\frac{k_{n}}{n^{2}}\cdot n\cdot\frac{n}{k_{n}}\cdot r_{n}^{-\gamma}\prec 1\quad\text{and}\quad[P1.2]\precsim\frac{k_{n}}{n^{2}}\cdot n\cdot\frac{n}{k_{n}}\cdot r_{n}^{-\gamma}\prec 1.$
@@
(SA.6.31)$\precsim\frac{k_{n}^{2}}{n^{4}}\sum_{i\in\mathcal{N}_{n}}\sum_{j\in\Lambda_{i}}\lvert\Xi_{ij}\rvert$ (SA.6.32)
@@
(SA.6.32)$\precsim\frac{k_{n}^{2}}{n^{4}}\cdot n\cdot\frac{n}{k_{n}}\cdot\frac{n^{2}}{k_{n}^{2}}\precsim k_{n}^{-1}\prec 1.$
@@
Then the claimed result for $\hat{\sigma}^{2}(2)$ follows from steps 1, 2, and 4 by replacing, for all $i,j$, every occurrence of “$A_{ij}(1)$” and “$\Lambda_{i}$” with “$A_{ij}(2)$” and “$C_{c(i)}$”, respectively. Write
@@
$\bigg{|}\frac{k_{n}}{n^{2}}\sum_{i\in\mathcal{N}_{n}}\sum_{j\in\Lambda_{i}\backslash C_{c(i)}}\text{Cov}({\bf E}[Z_{i}\mid\mathcal{F}_{i}],{\bf E}[Z_{j}\mid\mathcal{F}_{j}])\bigg{|}\\ \leq\frac{k_{n}}{n}\max_{i}\lvert\mathcal{N}(i,3U_{n})\backslash\mathcal{N}(i,L_{n})\rvert\precsim\frac{k_{n}\xi_{n}}{n}\left(\frac{n}{k_{n}\xi_{n}}\right)^{(d-1)/d}\prec 1.$

### v5

@@ 2.3 Causal Estimands
Because we allow for cross-cluster interference, we need to redefine conventional estimands in a manner free of this source of bias. To this end, define for $t\in\{0,1\}$ the $p_{t}$-counterfactual design, which sets $k=1$ and $q=t$ in Assumption 4. This design groups the population into a single cluster and assigns treatment with probability $p_{t}$. Throughout the paper, let $\bf E$$\cdot$] denote the expectation taken with respect to the observed design in Assumption 4 and ${\bf E}^{*}_{p_{t}}[\cdot]$ the expectation taken with respect to the $p_{t}$-counterfactual design.
@@ 5.1 Empirical Examples
Sur et al. (2009) conduct a CRT in an urban slum in India spanning about 1.2 by 0.7 km with 38k participants. To compute our suggested number of clusters (9), we need to select $\tilde{\gamma}$ and the unit of length. Suppose we choose the conservative bound $\tilde{\gamma}=d=2$, so that interference is assumed to decay like $r^{-2}$ or faster with each unit of length $r$. If we take 35m to be the unit of length, then for units $i,j,k$ such that $\rho(i,j)=35$m and $\rho(i,k)=70$m, the extent to which $k$’s treatment affects $i$ is less than $1/4$th ($2^{-\tilde{\gamma}}=0.25$) as small as the extent to which $j$’s affects $i$. This is with only a 35m difference in distance. For this unit of length, (9) yields $k=78$. In other words, these are the assumptions on interference that justify the authors’ choice of $k=80$, originally determined by a conventional power analysis based on a partial interference model.
@@ 6 Simulation Study
We conduct a simulation study to illustrate the finite-sample properties of our estimator and difference in means under our proposed design. We randomly draw unit locations from the square $(n\alpha_{n})^{1/2},(n\alpha_{n})^{1/2}]^{2}$ with $\alpha_{n}=0.8,0.7,0.6$, respectively. This corresponds to the infill-increasing case where the regional volume shrinks with the population size. We create clusters using $k$-medoids with $k$ given by (9) using the conservative choice $\tilde{\gamma}=d=2$ and set the assignment probabilities in Assumption 4 to $(q,p_{1},p_{0})=(0.7,0.5,0)$.
@@
Due to the choice of 5-5 in the spatial weights $w_{ij}$, Assumption 3 holds for $\gamma=3$ (Leung, 2022b, eq. (3)), which is a fairly slow rate of decay given that Assumption 3 requires $\gamma>2$.
@@
We present results for the indirect and total effects using 5000 simulation draws where within each draw, we redraw potential outcomes and recompute the design-based estimand. In Table 1, the “Spatial Interference” columns correspond to the outcome model described above, whereas the “Partial Interference” columns redefine $w_{ij}=0$ if $i,j$ lie in different clusters to eliminate cross-cluster interference. The “CI” rows report the coverage of 95-percent CIs using the indicated standard errors. For our estimator, the “SE” row corresponds to standard errors obtained from our variance estimator (4), while for difference in means, it corresponds to conventional cluster-robust SEs. The “SE” rows are the true superpopulation standard errors obtained by taking the standard deviation of the estimator across the simulation draws. As such, SE should be consistent for SE rather than conservative, while the CIs have asymptotically conservative coverage for the design-based estimand. Finally, row “% Excl” is the percentage of units that are not well surrounded.
@@ 7 Empirical Application
Egger et al. (2022) estimate the effects of the transfers on the population of eligible households using two main specifications described in their §3.2. Their “RF” (reduced form) specification is an OLS regression of an outcome on village- and sublocation-level treatment indicators and covariates. They report the coefficient on the village-level indicator, which corresponds to a model-assisted estimate of a village-level direct effect $\theta^{*}(1,0,2/3,2/3)$ among eligible households.
However, the authors note the potential for interference across sublocations (their quote in Example 2). For this reason, their preferred specification is the following “IV” (instrumental variables) regression. The main regressors are the amount of cash transferred per capita to the household’s village and the amount transferred to neighboring villages within a band of r2r-2 to $r$ km from the ego’s village for a range of $r$ values. The corresponding instruments are respectively an indicator for the ego’s village being treated and the share of eligible households in the band assigned to treatment. They use a BIC criterion to select the maximum range of $r$, which is 2 km. Using these estimates, the authors compute a model-assisted estimate of the total effect $\theta^{*}(1,0,2/3,0)$. This compares saturation levels of two-thirds and zero, which is nonparametrically unidentified since control villages have a saturation level of one-third. We instead report our estimates of the total effect $\theta^{*}(1,0,2/3,1/3)$.
@@
  • 653 villages (units), 84 sublocations (clusters). Standard errors are in parentheses. Column RF (IV) is the reduced form (IV) estimate of the direct (total) effect from Table I, column 1 (2) of Egger et al. (2022), $\hat{\theta}$ is our estimate, and $\hat{\theta}^{+}$ is difference in means. Our estimates use $r_{n}=1.6$, which results in 39.66 percent of units not being well surrounded.
@@ 8 Conclusion
    Initialize an arbitrary set of $k$ medoids $\mathcal{M}=\{m_{j}\}_{j=1}^{k}\subseteq\mathcal{N}_{n}$. Given any $\mathcal{M}$, the associated set of clusters is $\mathcal{C}_{n}=\{C_{j}\}_{j=1}^{k}$, where $C_{j}=\{i\in\mathcal{N}_{n}\colon m_{j}=\operatornamewithlimits{argmin}_{m\in\mathcal{M}}\rho(i,m)\}$.
@@
[…]
@@
It would also be useful to develop designs for nonparametrically estimating $\gamma$. A preliminary proposal is the following. Given $k$ clusters, say constructed using $k$-medoids, denote by $m_{j}$ the centroid of a cluster $C_{j}$. Randomize clusters to the following $T$ treatment arms. In the 0th arm, we assign all units to control. In the $t$th treatment arm for $t\geq 1$, we only assign units $i$ for which $\rho(i,m)\in(t,t+1]$ to treatment, where $m$ is the centroid of the cluster in question. That is, we treat only units in a ring at a certain distance from the centroid. Let $\hat{\theta}_{t}$ denote the difference-in-means estimate comparing average outcomes of units near centroids of clusters assigned to arm $t\geq 1$ with those assigned to arm 0. Then under Assumption 3, $\hat{\theta}_{t}$ should decay like $t^{-\gamma}$, producing a sort of “causal covariogram.” We may then regress $\log\hat{\theta}_{t}$ on $\log t$ to estimate $\gamma$.
@@
  • 653653 villages (units), $84$ sublocations (clusters). Standard errors are in parentheses. Column IV is the IV estimate of the overall effect from Table I, column 3 of Egger et al. (2022).
@@
[…]
@@
By Lemma SA.5.4, $\max_{i}\lvert\Lambda_{i}\rvert\precsim n/k_{n}$, so $P1]\precsim k_{n}^{-1}$. It remains to show that $P2]\precsim k_{n}^{-1}$. Intuitively, $[P2]$ should be well controlled since units not in $\Lambda_{i}$ are “far” from and therefore less correlated with $i$.
@@
[…]
$P2.1]=c_{n}^{\prime}\frac{k_{n}}{n^{2}}\frac{n}{k_{n}}\sum_{s=\lfloor 2r_{n}\rfloor}^{\lceil 4\bar{R}\rceil}C\xi_{n}s^{d-1}r_{n}^{-\gamma}\precsim\frac{\xi_{n}}{n}\bar{R}^{d}r_{n}^{-\gamma}$ (SA.6.10)
@@
[…]
@@
$P2.2]=c_{n}^{\prime}\frac{k_{n}}{n^{2}}\frac{n}{k_{n}}\sum_{s=\lfloor 4\bar{R}\rfloor}^{\infty}C\xi_{n}s^{d-1}(s/2-\bar{R})^{-\gamma}\\ \leq Cc_{n}^{\prime}\frac{\xi_{n}}{n}\sum_{s=\lfloor 4\bar{R}\rfloor}^{\infty}s^{d-1}(s/4)^{-\gamma}\precsim\frac{\xi_{n}}{n}\sum_{s=1}^{\infty}s^{d-1-\gamma}.$ (SA.6.13)
Combining (SA.6.9), (SA.6.10), (SA.6.12), and (SA.6.13) yields $P2]\precsim\xi_{n}/n\precsim k_{n}^{-1}$ since $k_{n}\xi_{n}/n\precsim 1$ by assumption.
@@
[…]
@@
$\tilde{\mathcal{B}}(C)=\big\{i\in C\colon\lvert\mathcal{N}(i,1)\cap(\{\xi_{n}^{-1/d}x\colon x\in\mathbb{Z}^{d}\}\backslash C)\rvert\geq 1\big\},$
@@
[…]
@@
[…]
@@
[…]
@@
$\lvert\tilde{\mathcal{B}}(\mathcal{N}_{n})\rvert\sim\xi_{n}\mathcal{R}_{n}^{d-1}\sim n(\xi_{n}/n)^{1/d}$ (SA.6.21)
Let $m_{j}$ denote the medoid of a cluster $C_{j}$. By Assumption 5(a), units in $\tilde{\mathcal{B}}(C_{j})$ are at least distance $\lfloor L_{n}\rfloor$ from $m_{j}$. Define the “contour set”
@@
$\lvert\tilde{\mathcal{B}}(C_{j})\rvert\geq\lvert\tilde{J}(\lfloor L_{n}\rfloor,C_{j})\rvert\sim\xi_{n}(n/(k_{n}\xi_{n}))^{(d-1)/d}$
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
$\begin{aligned} \frac{k_{n}}{n^{2}}\sum_{i\in\mathcal{N}_{n}}\sum_{j\in\mathcal{N}_{n}}(Z_{i}-{\bf E}[Z_{i}])&(Z_{j}-{\bf E}[Z_{j}])A_{ij}(1)\\ &-\frac{k_{n}}{n^{2}}\sum_{i\in\mathcal{N}_{n}}\sum_{j\in\mathcal{N}_{n}}\text{Cov}({\bf E}[Z_{i}\mid\mathcal{F}_{i}],{\bf E}[Z_{j}\mid\mathcal{F}_{j}])A_{ij}(1)\end{aligned}$
$\begin{aligned} =\frac{k_{n}}{n^{2}}\sum_{i\in\mathcal{N}_{n}}\sum_{j\in\mathcal{N}_{n}}(Z_{i}Z_{j}-{\bf E}[{\bf E}[Z_{i}\mid\mathcal{F}_{i}]&{\bf E}[Z_{j}\mid\,\mathcal{F}_{j}]])A_{ij}(1)\\ &-2\frac{k_{n}}{n^{2}}\sum_{i\in\mathcal{N}_{n}}(Z_{i}-{\bf E}[Z_{i}])\sum_{j\in\mathcal{N}_{n}}{\bf E}[Z_{j}]A_{ij}(1).\end{aligned}$
@@
As previously argued, $\max_{i}\sum_{j\in\mathcal{N}_{n}}{\bf E}[Z_{j}]A_{ij}(1)\precsim n/k_{n}$, and $\lvert n^{-1}\sum_{i\in\mathcal{N}_{n}}(Z_{i}-{\bf E}[Z_{i}])\rvert\prec 1$ by the proof of Theorem 2, so $P2]\prec 1$, while
$P1]=\frac{k_{n}}{n^{2}}\sum_{i\in\mathcal{N}_{n}}\sum_{j\in\mathcal{N}_{n}}(Z_{i}-{\bf E}[Z_{i}\mid\mathcal{F}_{i}])(Z_{j}-{\bf E}[Z_{j}\mid\mathcal{F}_{j}])A_{ij}(1)\\ +2\frac{k_{n}}{n^{2}}\sum_{i\in\mathcal{N}_{n}}\sum_{j\in\mathcal{N}_{n}}{\bf E}[Z_{i}\mid\mathcal{F}_{i}](Z_{j}-{\bf E}[Z_{j}\mid\mathcal{F}_{j}])A_{ij}(1)\\ +\frac{k_{n}}{n^{2}}\sum_{i\in\mathcal{N}_{n}}\sum_{j\in\mathcal{N}_{n}}({\bf E}[Z_{i}\mid\mathcal{F}_{i}]{\bf E}[Z_{j}\mid\mathcal{F}_{j}]-{\bf E}[{\bf E}[Z_{i}\mid\mathcal{F}_{i}]{\bf E}[Z_{j}\mid\mathcal{F}_{j}]])A_{ij}(1)\\ \equiv[P1.1]+[P1.2]+[P1.3].$
@@
$P1.1]\precsim\frac{k_{n}}{n^{2}}\cdot n\cdot\frac{n}{k_{n}}\cdot r_{n}^{-\gamma}\prec 1\quad\text{and}\quad[P1.2]\precsim\frac{k_{n}}{n^{2}}\cdot n\cdot\frac{n}{k_{n}}\cdot r_{n}^{-\gamma}\prec 1.$
@@
[…]
@@
[…]
@@
Then the claimed result for $\hat{\sigma}^{2}(2)$ follows from steps 1, 2, and 4 by replacing, for all $i,j$, every occurrence of $A_{ij}(1)$” and $\Lambda_{i}$” with $A_{ij}(2)$” and $C_{c(i)}$”, respectively. Write
@@
[…]